OpenStudy (anonymous):

six times the reciprocal of a number equals 3 times the reciprocal of 10?

OpenStudy (anonymous):

@texaschic101 can you help?

Mehek (mehek14):

Does it look like this: \(\tt{6*\dfrac{1}{n}=3*\dfrac{1}{10}}\)

OpenStudy (anonymous):

No,it is a word problem and I'm trying to write it down and find out what the number is

Mehek (mehek14):

Well I'm guessing it should look like this Does it look like this: \(\tt{6*\dfrac{1}{n}=3*\dfrac{1}{10}}\)

Mehek (mehek14):

and then you can solve for n

OpenStudy (anonymous):

20

Mehek (mehek14):

yes n = 20

OpenStudy (texaschic101):

yes